# Warm Inflation with Barrow Holographic Dark Energy

### v1

@@ 3.1 Warm Inflation in high dissipative regime
In this subsection, we assume that inflation occurs in a high dissipative regime i.e. $Q>>1$. Now we impose the condition on Eq(3.11) and get,
$\epsilon_{1}=\left[H^{-\Delta}\left\{\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}-1\right\}\right]$ (3.15)
@@
$\epsilon_{2}=\frac{\left[(-\Delta)H^{-(\Delta+1)}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}+H^{-\Delta}\left(1-\frac{\Delta}{2}\right)\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{\frac{-\Delta}{2}}\left(\dot{\alpha}+\dot{\beta}\frac{\dot{H}}{H^{2}}+\beta\left(\frac{H\ddot{H}-2\dot{H}^{2}}{H^{3}}\right)\right)\right]}{\left[H^{-\Delta}\left\{\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}-1\right\}\right]H}$ (3.16)
In our work we assume that the parameters $\alpha$ and $\beta$ are to be of the form $\alpha=\alpha_{0}H^{\gamma}$ and $\beta=\beta_{0}H^{\delta}$, where $\alpha_{0},\beta_{0},\gamma,\delta$ are constants. Imposing the condition $Q>>1$ and assuming $\alpha=\alpha_{0}H^{\gamma}$ and $\beta=\beta_{0}H^{\delta}$ Eq.(3.7) reads as
$\left(\alpha_{0}H^{\gamma+2}+\beta_{0}H^{\delta}\dot{H}\right)^{1-\frac{\Delta}{2}}-H^{2}=\frac{1}{2}\dot{H}$ (3.17)
Truncating the higher power of $H$ from Taylor series expansion of $(\alpha_{0}H^{\gamma+2}+\beta_{0}H^{\delta}\dot{H})^{1-\frac{\Delta}{2}}$ and using the relation in Eq.(2.8) we obtain $H$ as a function of the e-folding number $N$. Therefore we get
$H^{\delta-\gamma}={\alpha_{0}}^{1-\frac{\Delta}{2}}+C_{1}e^{\frac{-2(1-p)N}{A}}$ (3.18)
where $A=1-2(1-\frac{\Delta}{2})\beta_{0}{\alpha_{0}}^{-\frac{\Delta}{2}}$ and $p=\gamma+\delta+1$. For simplicity we assume that $\delta=\frac{\Delta\gamma}{2}+\Delta$. Now the slow roll parameters $\epsilon_{1}$, $\epsilon_{2}$ and $\beta_{in}$ can be obtained in terms of the e-folding number $N$. The reconstructed slow-roll parameters in terms of $N$ is given by
$\epsilon_{1}=-2\left(-1+P^{\Delta}\left(\alpha_{0}P^{-\gamma}-K\right)^{1-\frac{\Delta}{2}}\right)$ (3.19)
The second slow roll parameter in terms of $N$ is given by
$\epsilon_{2}=\frac{\gamma_{\epsilon 1}}{\gamma_{\epsilon 2}}$ (3.20)
@@
$\gamma_{\epsilon 1}=\left[]C_{1}e^{D}{\alpha_{0}}^{4}P^{\Delta}(\alpha_{0}P^{-\gamma}-K)^{-\frac{\Delta}{2}}\left({\alpha_{0}}^{2+\frac{\Delta}{2}}P^{-\gamma}(-2\gamma+\Delta\gamma+2\Delta)+\right.\right.$
$\left.\left.C_{1}e^{D}{\alpha_{0}}^{1+\Delta}P^{-\gamma}(-2\gamma+\Delta\gamma+2\Delta)+{\alpha_{0}}^{2}P^{-\gamma}\beta_{0}(\Delta-2)(-2\gamma+\Delta\gamma+2\Delta)-\right.\right.$
$\left.\left.(-2\gamma+\Delta\gamma+2\Delta)\beta_{0}P^{-\delta}2C_{1}e^{D}{\alpha_{0}}^{4}+{\alpha_{0}}^{1+\frac{\Delta}{2}}\beta_{0}(\Delta-2)\left(2P^{-\delta}(\gamma-\delta)+C_{1}e^{D}(-2\gamma+\Delta\gamma+2\Delta)\right)\right)\right]$ (3.21)
@@
$\beta_{in}=\frac{\gamma_{\beta_{1}}}{\gamma_{\beta_{1}}}$ (3.23)
@@
$\gamma_{\beta_{1}}=\left[C_{1}e^{D}m~{}\alpha_{0}P^{-2}\left(-4(\alpha_{0}+C_{1}e^{D}{\alpha_{0}}^{\frac{\Delta}{2}})({\alpha_{0}}^{\frac{\Delta}{2}}+\beta_{0}(\Delta-2))-\right.\right.$
$\left.\left.\left(P^{-2}(\alpha_{0}P^{-\gamma}-K)\right)^{-\frac{\Delta}{2}}\left({\alpha_{0}}^{2+\frac{\Delta}{2}}P^{-\gamma}(2+\gamma)+C_{1}e^{D}{\alpha_{0}}^{1+\Delta}P^{-\gamma}(2+\gamma)-2C_{1}e^{D}P^{-\delta}\beta_{0}(2+\delta)\right.\right.\right.$
$\left.\left.\left.+\left(2P^{-\delta}\beta_{0}(\delta-\gamma)+C_{1}e^{D}P^{-\gamma}(2+\gamma)(\Delta-2)\right)+{\alpha_{0}}^{2}P^{-\gamma}\beta_{0}(2+\gamma)(\Delta-2)\right)(\Delta-2)\right)\right]$ (3.24)
@@
Moreover the tensor-to-scalar ratio (from Eq.(2.12)) becomes,
$r=-\frac{\gamma_{r1}}{\gamma_{r2}}.$ (3.26)
@@
$\gamma_{r1}=\left[428.117P^{-1}(-1+P^{\Delta}(\alpha_{0}P^{-\gamma}-K)^{1-\frac{\Delta}{2}})\right].$ (3.27)
@@
$\gamma_{r2}=\left[\left(3^{-1+\frac{m}{4}}P\left(\frac{{M_{P}}^{2}}{\sigma_{r}}\left(P^{-2}-P^{-2}(\alpha_{0}P^{-\gamma}-K)\right)^{1-\frac{\Delta}{2}}\right)^{\frac{m}{4}}\right)^{2.8}\times\right.$
$\left.~{}~{}~{}~{}~{}~{}~{}~{}\left(\frac{{M_{P}}^{2}}{\sigma_{r}}(P^{-2}-(P^{-}2(\alpha_{0}P^{-\gamma}-K))^{1-\frac{\Delta}{2}})\right)^{\frac{1}{4}}\right]$ (3.28)
Using Eq.(2.11) the scalar spectral index is given by
$n_{s}=\frac{\gamma_{n_{1}}}{\gamma_{n_{2}}}$ (3.29)
@@
In the above equations, we have used $P$, $D$, $K$ whose values are
$P=\left(C_{1}e^{D}+{\alpha_{0}}^{1-\frac{\Delta}{2}}\right)^{\frac{1}{\gamma-\delta}}$ (3.32)
$K=\frac{2C_{1}{\alpha_{0}}^{\Delta}P^{-\delta}\beta_{0}}{\left(e^{D}\alpha_{0}+C_{1}{\alpha_{0}}^{\frac{\Delta}{2}}\right)\left({\alpha_{0}}^{\frac{\Delta}{2}}+\beta_{0}(\Delta-2)\right)}$ (3.33)
$D=\frac{2N(\delta-\gamma){\alpha_{0}}^{\frac{\Delta}{2}}}{{\alpha_{0}}^{\frac{\Delta}{2}}+\beta_{0}(\Delta-2)}$ (3.34)
3.2 Model comparison with observational data
@@
In this section, we compare the results with observational data to validate the model. In the previous section we have already expressed $\epsilon_{1}$, $\epsilon_{2}$, $r$, $\beta_{in}$, $n_{s}$ in terms of e-folding number $N$. Warm inflation occurs when slow roll parameters meet the condition $\left|\epsilon_{1}\right|<<1$,  $\left|\epsilon_{2}\right|<<1$, and  $\left|\beta\right|<<1$. From Fig.1, Fig.2, and Fig.3 it is apparent that all the three slow roll parameters satisfy the conditions for warm inflation.
Figure 1: Plot of the first slow-roll parameter $\epsilon_{1}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are considered as $C_{1}=2.05$$\gamma=0.04$$\alpha_{0}=2.36$$\beta_{0}=2.21$.
@@
Figure 2: Plot of the slow-roll parameter $\epsilon_{2}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are considered as $C_{1}=200.05$$\gamma=5.04$$\alpha_{0}=0.036$$\beta_{0}=0.21$.
@@
Figure 3: Plot of the slow-roll parameter $\beta_{in}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. Other parameters are considered as $C_{1}=0.2$$\gamma=2.04$$\alpha_{0}=0.036$$\beta_{0}=0.21$.
Fig.4 illustrates spectral index $n_{s}$ versus e-folding number $N$ for different values of the Barrow parameter $\Delta$. The latest observational data states that $n_{s}$ lies in the range $0.9642\pm 0.0042$ [53]. In our model, the spectral index nearly lies between the above-mentioned range. We also see that for a greater value of $\Delta$, we get a greater value of $n_{s}$. In Fig.5 we have plotted the tensor-to-scalar ratio $r$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. From the latest observational data, it is known that the upper limit of $r$ is $r<0.064$. From the plot, we see that for our model $r$ lies in this admissible range. It is also seen that generally, for a higher value of $\Delta$, we get a higher value of $r$.
Figure 4: Plot of the scalar spectral index $n_{s}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The parameters considered here are $\sigma_{r}=1$, $\gamma=0.002$, $\alpha_{0}=0.03$, $\beta_{0}=0.21$, $m=2.08$.
@@
Figure 5: Plot of the tensor-to-scalar ratio $r$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are considered as $M_{p}=1$, $\sigma_{r}=1$, $c_{1}=0.2$, $\gamma=2.04$, $B_{t}=4.5$, $\alpha_{0}=0.036$, $\beta_{0}=0.21$, $m=0.025$.
Fig.6 illustrates the behavior of energy density of BHDE for different values of the Barrow parameter. It is apparent that energy density was high during the initial time and decreased over time which is quite expected in an expanding scenario. Moreover, it shows that energy was transmitted from BHDE to radiation during the inflation. We do not get a clear comparison depending on the values of $\Delta$, since there is a cross-over between the trajectories. In Fig.7 the ratio of the holographic energy density and radiation density is plotted. At the initial time, the ratio was high however, at the end of the inflation the ratio decreases as the ratios come close to each other.
Figure 6: Plot of the holographic energy density $\rho_{BHDE}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The parameters chosen here are $\sigma=1$, $C_{1}=10^{5}$, $\gamma=3.04$, $\alpha_{0}=800.36$, $\beta_{0}=90.21$.
@@
Figure 7: Plot of $\frac{\rho_{BHDE}}{\rho_{r}}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are considered as $M_{p}=1$, $\sigma_{r}=1$, $C_{1}=20\times 10^{5}$, $\gamma=1.10574$, , $\alpha_{0}=1.0836$, $\beta_{0}=0.0921$.
@@
Finally, we have observed the behavior of $T$ and $H$ in Fig.8. In Fig.8 it is seen that $\frac{T}{H}$ is very much greater than 1 i.e in the presence of a thermal bath when $T>H$ the quantum fluctuations of the fields are dominated by the thermal fluctuations. So our model perfectly supports a warm inflationary scenario.

### v2

@@ 3.1 Warm Inflation in high dissipative regime
In this subsection, we assume that inflation occurs in a high dissipative regime i.e. $Q>>1$. Now we impose the condition on Eq. (3.11) and get,
$\epsilon_{1}=-2\left[H^{-\Delta}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}-1\right].$ (3.15)
@@
$\epsilon_{2}=\frac{\left[(-\Delta)H^{-(\Delta+1)}\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}+H^{-\Delta}\left(1-\frac{\Delta}{2}\right)\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{\frac{-\Delta}{2}}\left(\dot{\alpha}+\dot{\beta}\frac{\dot{H}}{H^{2}}+\beta\left(\frac{H\ddot{H}-2\dot{H}^{2}}{H^{3}}\right)\right)\right]}{\left[H^{-\Delta}\left\{\left(\alpha+\beta\frac{\dot{H}}{H^{2}}\right)^{1-\frac{\Delta}{2}}-1\right\}\right]H}.$ (3.16)
In our work, we assume that the parameters $\alpha$ and $\beta$ are to be of the form $\alpha=\alpha_{0}H^{\gamma}$ and $\beta=\beta_{0}H^{\delta}$, where $\alpha_{0},\beta_{0},\gamma,\delta$ are constants. Imposing the condition $Q>>1$ and assuming $\alpha=\alpha_{0}H^{\gamma}$ and $\beta=\beta_{0}H^{\delta}$ Eq.(3.7) reads as,
$\left(\alpha_{0}H^{\gamma+2}+\beta_{0}H^{\delta}\dot{H}\right)^{1-\frac{\Delta}{2}}-H^{2}=\frac{1}{2}\dot{H}.$ (3.17)
Truncating the higher power of $H$ from Taylor series expansion of $(\alpha_{0}H^{\gamma+2}+\beta_{0}H^{\delta}\dot{H})^{1-\frac{\Delta}{2}}$ and using the relation in Eq.(2.9) we obtain $H$ as a function of the e-folding number $N$. Therefore we get,
$H^{\delta-\gamma}={\alpha_{0}}^{1-\frac{\Delta}{2}}+C_{1}e^{\frac{-2(1-p)N}{A}},$ (3.18)
where $A=1-2(1-\frac{\Delta}{2})\beta_{0}{\alpha_{0}}^{-\frac{\Delta}{2}}$ and $p=\gamma+\delta+1$. For simplicity we assume that $\delta=\frac{\Delta\gamma}{2}+\Delta$. Now the slow roll parameters $\epsilon_{1}$, $\epsilon_{2}$, and $\beta_{in}$ can be obtained in terms of the e-folding number $N$. The reconstructed slow-roll parameters in terms of $N$ is given by
$\epsilon_{1}=-2\left(-1+P^{\Delta}\left(\alpha_{0}P^{-\gamma}-K\right)^{1-\frac{\Delta}{2}}\right).$ (3.19)
The second slow roll parameter in terms of $N$ is given by,
$\epsilon_{2}=\frac{\gamma_{\epsilon 1}}{\gamma_{\epsilon 2}},$ (3.20)
@@
$\gamma_{\epsilon 1}$ $=$ $C_{1}e^{D}{\alpha_{0}}^{4}P^{\Delta}(\alpha_{0}P^{-\gamma}-K)^{-\frac{\Delta}{2}}\left({\alpha_{0}}^{1+\frac{\Delta}{2}}\beta_{0}(\Delta-2)\left(2P^{-\delta}(\gamma-\delta)+C_{1}e^{D}(-2\gamma+\Delta\gamma+2\Delta)\right)\right)$ (3.21)
$+$ $C_{1}e^{D}{\alpha_{0}}^{4}P^{\Delta}(\alpha_{0}P^{-\gamma}-K)^{-\frac{\Delta}{2}}\left(C_{1}e^{D}{\alpha_{0}}^{1+\Delta}P^{-\gamma}(-2\gamma+\Delta\gamma+2\Delta)\right)$
$+$ $C_{1}e^{D}{\alpha_{0}}^{4}P^{\Delta}(\alpha_{0}P^{-\gamma}-K)^{-\frac{\Delta}{2}}\left({\alpha_{0}}^{2}P^{-\gamma}\beta_{0}(\Delta-2)(-2\gamma+\Delta\gamma+2\Delta)\right)$
-- $C_{1}e^{D}{\alpha_{0}}^{4}P^{\Delta}(\alpha_{0}P^{-\gamma}-K)^{-\frac{\Delta}{2}}\left((-2\gamma+\Delta\gamma+2\Delta)\beta_{0}P^{-\delta}2C_{1}e^{D}{\alpha_{0}}^{4}\right)$
$+$ $C_{1}e^{D}{\alpha_{0}}^{4}P^{\Delta}(\alpha_{0}P^{-\gamma}-K)^{-\frac{\Delta}{2}}\left({\alpha_{0}}^{2+\frac{\Delta}{2}}P^{-\gamma}(-2\gamma+\Delta\gamma+2\Delta)\right)$
@@
$\beta_{in}=\frac{\gamma_{\beta_{1}}}{\gamma_{\beta_{2}}}$ (3.23)
@@
$\gamma_{\beta_{1}}=C_{1}e^{D}m~{}\alpha_{0}P^{-2}\left(X\right),$ (3.24)
with
$X$ $=$ $-4(\alpha_{0}+C_{1}e^{D}{\alpha_{0}}^{\frac{\Delta}{2}})({\alpha_{0}}^{\frac{\Delta}{2}}+\beta_{0}(\Delta-2))$
-- $\left(P^{-2}(\alpha_{0}P^{-\gamma}-K)\right)^{-\frac{\Delta}{2}}\left(Y\right)(\Delta-2),$
and
$Y$ $=$ ${\alpha_{0}}^{2+\frac{\Delta}{2}}P^{-\gamma}(2+\gamma)+C_{1}e^{D}{\alpha_{0}}^{1+\Delta}P^{-\gamma}(2+\gamma)-2C_{1}e^{D}P^{-\delta}\beta_{0}(2+\delta)$
$+$ $\left(2P^{-\delta}\beta_{0}(\delta-\gamma)+C_{1}e^{D}P^{-\gamma}(2+\gamma)(\Delta-2)\right)+{\alpha_{0}}^{2}P^{-\gamma}\beta_{0}(2+\gamma)(\Delta-2).$
@@
Now we will deduce the scalar spectral index and the tensor-to-scalar ratio. Since we are in HDR the dissipative parameter $Q$ is bigger than one. Therefore, the function $G(Q)$ in Eq.(2.10) is approximated as $G(Q)\approx 0.0185Q^{2.315}$ [74]. Using Eq.(2.12) we obtain the scalar spectral index as [74],
$n_{s}=1+1.815\epsilon_{1}-\epsilon_{2}+3.815\beta.$ (3.26)
Again the tensor-to-scalar ratio is obtained by computing Eq.(2.10) and Eq.(2.13) as [74],
$r=16\epsilon_{1}\left(\sqrt{3}\pi\frac{T}{H}0.0185Q^{2.815}\right)^{-1}.$ (3.27)
Therefore the tensor-to-scalar ratio (from Eq.(3.27)) becomes,
$r=-\frac{\gamma_{r1}}{\gamma_{r2}},$ (3.28)
@@
$\gamma_{r1}=\left[428.117P^{-1}(-1+P^{\Delta}(\alpha_{0}P^{-\gamma}-K)^{1-\frac{\Delta}{2}})\right],$ (3.29)
@@
$\gamma_{r2}$ $=$ $\left(3^{-1+\frac{m}{4}}P\left(\frac{{M_{P}}^{2}}{\sigma_{r}}\left(P^{-2}-P^{-2}(\alpha_{0}P^{-\gamma}-K)\right)^{1-\frac{\Delta}{2}}\right)^{\frac{m}{4}}\right)^{2.8}$ (3.30)
$\times$ $\left(\frac{{M_{P}}^{2}}{\sigma_{r}}(P^{-2}-(P^{-2}(\alpha_{0}P^{-\gamma}-K))^{1-\frac{\Delta}{2}})\right)^{\frac{1}{4}}$
Using Eq.(3.26) the scalar spectral index is given by,
$n_{s}=\frac{\gamma_{n_{1}}}{\gamma_{n_{2}}},$ (3.31)
@@
In the above equations, we have used $P$, $D$, and $K$ whose values are
$P=\left(C_{1}e^{D}+{\alpha_{0}}^{1-\frac{\Delta}{2}}\right)^{\frac{1}{\gamma-\delta}},$ (3.34)
$K=\frac{2C_{1}{\alpha_{0}}^{\Delta}P^{-\delta}\beta_{0}}{\left(e^{D}\alpha_{0}+C_{1}{\alpha_{0}}^{\frac{\Delta}{2}}\right)\left({\alpha_{0}}^{\frac{\Delta}{2}}+\beta_{0}(\Delta-2)\right)},$ (3.35)
$D=\frac{2N(\delta-\gamma){\alpha_{0}}^{\frac{\Delta}{2}}}{{\alpha_{0}}^{\frac{\Delta}{2}}+\beta_{0}(\Delta-2)},$ (3.36)
In this high dissipative regime analysis, we have successfully derived the key parameters characterizing the warm inflationary dynamics driven by Barrow holographic dark energy. The slow-roll parameters $\epsilon_{1}$, $\epsilon_{2}$ and the warm inflation parameter $\beta_{in}$ have been expressed explicitly in terms of the e-folding number $N$. These expressions incorporate both the effects of the holographic nature of dark energy through the Barrow parameter $\Delta$ and the dissipative dynamics through the coefficient $\Gamma$. This theoretical framework sets the stage for a detailed comparison with observational constraints, particularly from cosmic microwave background measurements, which we shall explore in the following subsection.
An important result of our analysis is that the reconstructed parameters naturally encode the fractal characteristics inherited from the Barrow entropy formulation while maintaining consistency with the requirements of warm inflation. The high dissipative regime ($Q\gg 1$) ensures efficient energy transfer from the holographic dark energy to radiation, which is crucial for sustaining the warm inflationary phase. This efficient dissipation mechanism, combined with the modified holographic entropy-area scaling, provides a novel perspective on the early universe dynamics.
3.2 Results
@@
In this section, we discuss the results and compare the results with observed values to validate the model. In the previous section we have already expressed $\epsilon_{1}$, $\epsilon_{2}$, $r$, $\beta_{in}$, and $n_{s}$ in terms of e-folding number $N$. Warm inflation occurs when slow roll parameters meet the condition $\left|\epsilon_{1}\right|<<1$,  $\left|\epsilon_{2}\right|<<1$, and  $\left|\beta\right|<<1$. From Fig.(1), Fig.(2), and Fig.(3) it is apparent that all three slow roll parameters satisfy the conditions for warm inflation. The most important idea is that, if inflation stops i.e at $N=0$, the first slow-roll parameter $\epsilon_{1}$ tends to one that indicates the end of the accelerated expansion of the universe [74]. We have computed $\epsilon_{1}$ at $N=0$ by choosing suitable values of the parameters and we have shown that $\epsilon_{1}\approx 1$ and $\epsilon_{1}<<1$ in the region of 40-70 e-folding. Tables 1, 2, and 3 can give more insight into this where some numerical results have been presented.
@@
Figure 1: Plot of the slow-roll parameter $\epsilon_{1}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are considered as $C_{1}=1.2$$\gamma=1.02$$\alpha_{0}=0.06$$\beta_{0}=0.01$.
@@
Figure 2: Plot of the slow-roll parameter $\epsilon_{2}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are considered as $C_{1}=1.2$$\gamma=1.02$$\alpha_{0}=0.06$$\beta_{0}=0.01$.
@@
Figure 3: Plot of the slow-roll parameter $\beta_{in}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. Other parameters are considered as $C_{1}=1.2$$\gamma=1.02$$\alpha_{0}=0.06$$\beta_{0}=0.01$.
Fig.(4) illustrates spectral index $n_{s}$ versus e-folding number $N$ for different values of the Barrow parameter $\Delta$. The latest observational data states that $n_{s}$ lies in the range $0.9642\pm 0.0042$ [79]. In our model, the spectral index nearly lies between the above-mentioned range. In Fig.(5) we have plotted the tensor-to-scalar ratio $r$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. From the latest observational data, it is known that the upper limit of $r$ is $r<0.064$. From the plot, we see that for our model $r$ lies in this admissible range. It is also seen that generally, for a higher value of $\Delta$, we get a higher value of $r$.
Table 1: The table shows numerical values of $\epsilon_{1}$ for different e-folding numbers for $\Delta=0.65$
N 0 10 20 30 40 50 60 70
$C_{1}$ 1.02 1.02 1.02 1.02 1.02 1.02 1.02 1.02
$\gamma$ 0.84 0.84 0.84 0.84 0.84 0.84 0.84 0.84
$\alpha_{0}$ 1.06 1.06 1.06 1.06 1.06 1.06 1.06 1.06
$\beta_{0}$ 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01
$\epsilon_{1}$ 1.0035 0.3126 0.066 0.012 0.0023 0.00044 0.000082 0.000015
Table 2: The table shows numerical values of $\epsilon_{1}$ for different e-folding numbers for $\Delta=0.67$.
N 0 10 20 30 40 50 60 70
$C_{1}$ 1.02 1.02 1.02 1.02 1.02 1.02 1.02 1.02
$\gamma$ 0.84 0.84 0.84 0.84 0.84 0.84 0.84 0.84
$\alpha_{0}$ 1.06 1.06 1.06 1.06 1.06 1.06 1.06 1.06
$\beta_{0}$ 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01
$\epsilon_{1}$ 1.0036 0.1886 0.02153 0.00227 0.00023 0.000024 2.60×$\times 10^{-6}$ 2.72×$\times 10^{-6}$
Table 3: The table shows numerical values of $\epsilon_{1}$ for different e-folding numbers for $\Delta=0.69$.
N 0 10 20 30 40 50 60 70
$C_{1}$ 1.02 1.02 1.02 1.02 1.02 1.02 1.02 1.02
$\gamma$ 0.84 0.84 0.84 0.84 0.84 0.84 0.84 0.84
$\alpha_{0}$ 1.06 1.06 1.06 1.06 1.06 1.06 1.06 1.06
$\beta_{0}$ 0.01 0.01 0.01 0.01 0.01 0.01 0.01 0.01
$\epsilon_{1}$ 1.0037 0.11070 0.0068 0.00040 0.000023 1.40×$\times 10^{-6}$ 8.28×$\times 10^{-8}$ 4.87×$\times 10^{-9}$
@@
Figure 4: Plot of the scalar spectral index $n_{s}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The parameters considered here are $\sigma_{r}=1$, $\gamma=1.02$, $\alpha_{0}=0.06$, $\beta_{0}=0.01$, $m=1$, $C_{1}=1.2$.
Figure 5: Plot of the tensor-to-scalar ratio $r$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are considered as $\sigma_{r}=1$, $C_{1}=1.2$, $\gamma=1.02$, $B_{t}=0.5$, $\alpha_{0}=0.06$, $\beta_{0}=0.01$, $m=1$.
Figure 6: Plot of the tensor-to-scalar ratio $r$ versus the scalar spectral index $n_{s}$ for different values of e-folding number $N$. The other parameters are considered as $M_{p}=1$, $\sigma_{r}=1$, $C_{1}=1.2$, $\gamma=1.02$, $B_{t}=0.5$, $\beta_{0}=0.01$, $m=1$$\Delta=0.048$.
Fig.(6) illustrates tensor-to-scalar ratio $r$ versus the scalar spectral index $n_{s}$ for different values of e-folding number $N$. It is seen that with increasing $n_{s}$ there is a corresponding increase in $r$. Here the values of $n_{s}$ and $r$ correspond to the values at the time of horizon crossing. In Table 4 some numerical results are presented to get a better idea on these parameters. Fig.(7) illustrates the behavior of energy density of BHDE for different values of the Barrow parameter. At first, it has a high value, but it decreases with the evolution of time. Inflation has an energy density of roughly $10^{64}$ GeV. Some energy was transmitted from BHDE to radiation during the inflation. We do not get a clear comparison depending on the values of $\Delta$, since there is a cross-over between the trajectories. In Fig.(8) the ratio of the holographic energy density and radiation density is plotted against $N$. At the initial time, the ratio was high, however, at the end of the inflation the ratio decreased and the different trajectories came close to each other. The value of the ratio is much greater than $1$, at the initial time, which shows that the inflation was driven by BHDE and $\rho_{BHDE}$ dominates. As we reach the end of the inflation we see that the ratio gets smaller, which means that the two densities approach each other.
Figure 7: Plot of the holographic energy density $\rho_{BHDE}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are $\sigma_{r}=1$, $C_{1}=1.2$, $\gamma=1.02$, $\alpha_{0}=0.06$, $\beta_{0}=0.01$.
Figure 8: Plot of $\frac{\rho_{BHDE}}{\rho_{r}}$ against the e-folding number $N$ for different values of the Barrow parameter $\Delta$. The other parameters are considered as $M_{p}=1$, $\sigma_{r}=1$, $C_{1}=1.2$, $\gamma=1.02$, , $\alpha_{0}=0.06$, $\beta_{0}=0.01$.
@@
Finally, we have observed the behavior of $T$ and $H$ in Fig.(9) and dissipative parameter $Q$ in Fig.(10). In Fig.(9) it is seen that $\frac{T}{H}$ is very much greater than 1 i.e in the presence of a thermal bath when $T>H$ the quantum fluctuations of the fields are dominated by the thermal fluctuations. So our model perfectly supports a warm inflationary scenario. In Tables 5 and 6 some numerical values of $Q$ and $T/H$ have been presented to support the idea obtained from the plots.